OpenStudy (anonymous):

x+4 over x added to x over x+4. I think i over simplified the problem cause i cant get the right answer. Help please

jhonyy9 (jhonyy9):

x+4 x (x+4)^2 +x^2 ---- + ----- = ------------- x x+4 x(x+4)

OpenStudy (campbell_st):

you have \[\frac{x+4}{x}+\frac{x}{x+4}\] you need a common denominator so... multiply the denominators and then cross multiply the numerator with the other denominator \[\frac{(x+4)(x+4)}{x(x+4)} + \frac{ x \times x}{x(x+4)}\] gives \[\frac{(x+4)^2 + x^2}{x(x+4}\] \[\frac{x^2 + 8x + 16 +x^2}{x^2 + 4x} \] I'll leave you to finish it off

OpenStudy (campbell_st):

oops don't expand the denominator

jhonyy9 (jhonyy9):

x^2 +8x +16 +x^2 2x^2 +8x +16 ------------------ = -------------- x(x+4) x(x+4)

OpenStudy (anonymous):

WHy couldnt you just add a plus four to the denominator to make them x+4?

OpenStudy (campbell_st):

doesn't quite work like that... but I like the idea... it simplifies maths... a lot

OpenStudy (campbell_st):

the simplified version is \[\frac{2(x^2 + 4x + 8)}{x(x+4)}\]

OpenStudy (anonymous):

Okay, im trying to work it out without looking too much as your answers to see if i can figure it out. the fact that i cant just add the plus four still confuses me

OpenStudy (campbell_st):

as a simple example... what is 1/2 + 1/3 = 5/6 you need a common denominator 2 x 3 and cross multply the numerators... 1 x 2 + 1x 3 = 5